# Waveform proportionality and Taylor’s law in coupled Lorenz systems

### v1

@@ Abstract
Taylor’s law (TL) describes a power-law relationship between the mean and variance of a quantity, a phenomenon observed across diverse scientific disciplines. Despite its ubiquity, the underlying mechanisms responsible for TL are not yet fully elucidated. In particular, the frequent empirical observation of TL with an exponent 2 warrants further investigation. In a previous study [Phys. Rev. Lett. 134, 167202 (2025)], we hypothesized that synchronization contributes to the emergence of TL with an exponent 2. We validated this hypothesis using coupled oscillator models derived from a food chain model, the Rössler system, the Brusselator, and the Lorenz system. Our analytical and numerical results demonstrated that strong coupling leads to a form of synchronization wherein time series become proportional to each other, consequently resulting in TL with an exponent 2. Here, we extend our previous findings for the coupled Lorenz system and provide detailed calculations.
@@ I Introduction
When the exponent takes values greater than one, this power-law relationship is often termed giant (number/density) fluctuations, a phenomenon extensively studied in the field of active matter physics [5, 6, 7]. Empirical data on population abundance in ecosystems frequently yield TL exponents close to 2 [8, 9, 10]. However, theoretical investigations have demonstrated that the TL exponent can take arbitrary real values [11, 12, 13], stimulating interest in understanding the mechanisms underlying the frequent emergence of TL with an exponent 2 [14]. In this context, we posited that synchronization underlies the emergence of TL with an exponent 2 [15], as several studies have indicated that stronger correlations between time series tend to drive the TL exponent towards 2 [9, 16, 17]. Intriguingly, when the correlation becomes sufficiently strong such that the time series become proportional to each other, spatial TL with an exponent 2 emerges [18]. In Ref. [15], we investigated our hypothesis using coupled dynamical system models and discovered that strong coupling induces a type of synchronization where time series become mutually proportional, resulting in TL with an exponent 2. We term this synchronization type waveform proportionality (WP). Although synchronization characterized by proportional time series has been previously reported as projective synchronization [19] or as a specific instance of generalized synchronization [20], these observations were confined to particular classes of models. More recently, we demonstrated that this synchronization type can manifest in a broader range of models [15]. In Ref. [15], we employed the coupled food chain model [21, 22, 23], the coupled Rössler system [24, 25, 26, 27, 23, 28], the coupled Brusselator [29], and the coupled Lorenz system [30]. For the initial three models, we elucidated the mechanism of TL using a perturbative method and an averaging approximation. In contrast, the coupled Lorenz system necessitated a self-consistent approach instead of the averaging method. In this paper, we extend our previous results for the coupled Lorenz system and provide detailed calculations. We present results not only for the coupled Lorenz system [30] which is based on the original Lorenz system [31], but also for coupled systems constructed from the generalized Lorenz system [32, 33] and the hyperchaotic Lorenz system [34]. The generalized Lorenz system, an extension of the original Lorenz system incorporating the influence of a magnetic field [32], has been shown to exhibit intriguing intermittent dynamics [33]. The hyperchaotic Lorenz system [34], capable of possessing two positive Lyapunov exponents, displays more complex dynamics than the original Lorenz system [31] and is of particular interest for applications in secure communication and cryptography.
@@ II.3 Numerical simulations
where $\mathrm{CV}$ denotes the coefficient of variation. Namely, $\chi$ is the CV of the mean field of $z_{i}(t)$ normalized by the maximum CV among the individual $z_{i}(t)$. With this definition, $0\leq\chi\leq 1$, where $\chi\simeq 0$ indicates an unsynchronized state and $\chi\simeq 1$ signifies a well-synchronized state. The quantity $\zeta_{i}(t)$ represents the ratio between time series; it becomes approximately constant when WP holds. While the choice of $j$ is arbitrary, we select $j=50$, assuming the parameters $\mu_{i}$ satisfy $\mu_{1}<\mu_{2}<\cdots<\mu_{100}$. By arranging them in this manner, we can expect $z_{50}(t)\simeq z_{0}(t)$. Unless otherwise specified, $\mu_{i}$ is selected from a uniform distribution between $-1.5$ and $1.5$.
@@
Next, we examine the dependence of the TL parameters on the coupling strength. According to the analytical derivation, the numerical simulation results are expected to agree well with the theoretical predictions as $D_{x}$ increases. We verify this expectation using the coupled Lorenz system [30]; the results are shown in Fig. 2. Simulations are performed up to $t=3500$, and TL is computed using the time series from $t=3000$ to $t=3500$. The error bars represent the standard deviation of ten calculations with different initial conditions and realizations of $\rho_{i}$. Initial conditions for numerical simulations are randomly chosen from a uniform distribution between 0 and 1. Depending on the coupling strength, oscillation quenching may occur, rendering it impossible to define $\chi$. Thus, we judge that oscillation quenching occurs when $(1/N)\sum_{i=1}^{N}\langle\left(z_{i}-\langle z_{i}\rangle_{t}\right)^{2}\rangle_{t}$ falls below a certain threshold value; such cases are excluded. We can confirm that qualitatively the same results are obtained for several different threshold values. This procedure is applied to all systems described subsequently. Figure 2 illustrates that as the coupling strength increases, $\chi$ begins to rise, indicating the onset of synchronization, when $D_{x}$ becomes comparable to $\max\{\mu_{i}\}$. With a further increase in coupling strength, $\chi\simeq 1$ and $R^{2}\simeq 1$, and the theoretical predictions and numerical simulation results for both $\log\alpha_{\rm t,s}$ and $\beta_{\rm t,s}$ show excellent agreement.
Figure 2: Dependence of TL parameters on $D_{x}$ in the coupled Lorenz system. $D_{y}=D_{z}=0,N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (c), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
@@ III.1 Model
where $x$, $y$, and $z$ are the original variables of the Lorenz system, and $\sigma$, $\rho$, and $b$ are its original parameters. The variable $w$ and parameters $a$, $c$, $k$, and $l$ are newly introduced to extend the original Lorenz system. By appropriately choosing the parameters $a$, $c$, $k$, and $l$, we can recover the original/generalized/hyperchaotic Lorenz system; when $a=c=k=l=0$, the system reduces to the original Lorenz system (2); when $a=-\omega_{0}$, $c=\omega_{0}$, $k=0$, and $l=\sigma_{\rm m}$, it corresponds to the generalized Lorenz system (24); and when $a=1$, $c=0$, $k=1$, and $l=-\gamma$, it corresponds to the hyperchaotic Lorenz system (25). We refer to the system described by Eq. (26) as the extended Lorenz system. Following the formulation of the coupled Lorenz system used in Ref. [30], we construct the following coupled system for the extended Lorenz system:
@@ III.2 Analytical derivation
When $|\sigma(y_{0}-x_{0})|,|aw_{0}|\gg|\mu_{i}p|,|{\varepsilon}_{i}\sigma q|,|\mu_{i}\tau\sigma(y_{0}-x_{0})|,|\mu_{i}\tau aw_{0}|$, the following equations can be derived by following the same procedure as before.
@@ III.3 Numerical simulations
Figure 3: Dependence of TL parameters on $D_{x}$ in the coupled generalized Lorenz system. $a=-1,c=1,k=0,l=20$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (c), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Figure 4: Dependence of TL parameters on $D_{x}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (c), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
@@
Next, we investigate the effects of the other coupling coefficients, $D_{y},D_{z},$ and $D_{w}$, on the validity of the theory. Based on the derivation, we expect that the analytical framework to become invalid when $D_{y}$ or $D_{z}$ is non-negligible. In contrast, our theory predicts that increasing $D_{w}$ does not affect the results. To test these expectations, we perform numerical simulations using the coupled hyperchaotic Lorenz system (Figs. 6, 7, and 8).
@@
Figure 8: Dependence of TL parameters on $D_{w}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{x}=1000,D_{y}=D_{z}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
@@ IV Discussion
In this study, building upon our previous work [15], we investigated the emergence of WP and the associated TL with an exponent 2 in coupled Lorenz systems. In particular, we extended the results obtained in [15] by considering a coupled system constructed from the extended Lorenz system (26), which is newly introduced in this study. We verified our analytical derivation through numerical simulations using coupled systems constructed from the original, generalized, and hyperchaotic Lorenz systems. In the case of the four-variable model, we numerically observe that $|aw_{0}|$ is not always large and our assumption of large $|aw_{0}|$ in our perturbative approach might be violated. Nevertheless, the numerical simulation results are still in good agreement with the theoretical predictions. This observation suggests that the treatment of the $aw_{0}$ term in our perturbative approach requires further consideration.

### v2

@@ Abstract
Taylor’s law (TL), a power-law relationship between the mean and variance of a quantity, has been observed across diverse scientific disciplines. Despite its ubiquity, the underlying mechanisms responsible for TL are not yet fully elucidated. In particular, the frequent empirical observation of TL with an exponent 2 warrants further investigation. In a previous study [Phys. Rev. Lett. 134, 167202 (2025)], we hypothesized that synchronization contributes to the emergence of TL with an exponent 2. To validate this hypothesis, we employed coupled oscillator models, with each oscillator described by a distinct dynamical system: a food chain model, the Rössler system, the Brusselator, and the Lorenz system. Our analytical and numerical results demonstrated that strong coupling leads to a form of synchronization wherein time series become proportional to each other, consequently resulting in TL with an exponent 2. Here, we extend our previous findings for the coupled Lorenz system and provide detailed calculations. Our analytical and numerical results demonstrate that, under strong coupling, waveform proportionality and Taylor’s law with an exponent 2 emerge not only in the original Lorenz system but also in the generalized and hyperchaotic Lorenz systems.
@@ I Introduction
When the exponent takes values greater than one, this power-law relationship is often referred to as giant (number/density) fluctuations, a phenomenon extensively studied in the field of active matter physics [5, 6, 7]. Empirical data on population abundance in ecosystems frequently yield TL exponents close to 2 [8, 9, 10]. However, theoretical investigations have demonstrated that the TL exponent can take arbitrary real values [11, 12, 13], stimulating interest in understanding the mechanisms underlying the frequent emergence of TL with an exponent 2 [14]. In this context, we posited that synchronization underlies the emergence of TL with an exponent 2 [15], as several studies have indicated that stronger correlations between time series tend to drive the TL exponent towards 2 [9, 16, 17]. Intriguingly, when the correlation becomes sufficiently strong such that the time series become proportional to each other, spatial TL with an exponent 2 emerges [18]. In Ref. [15], we investigated our hypothesis using coupled dynamical system models and discovered that strong coupling induces a type of synchronization where time series become mutually proportional, resulting in TL with an exponent 2. We refer to this type of synchronization as waveform proportionality (WP). Although synchronization characterized by proportional time series has been previously reported as projective synchronization [19] or as a specific instance of generalized synchronization [20], these observations were confined to particular classes of models. More recently, we demonstrated that this type of synchronization can manifest in a broader range of models [15]. In Ref. [15], we employed the coupled food chain model [21, 22, 23], the coupled Rössler system [24, 25, 26, 27, 23, 28], the coupled Brusselator [29], and the coupled Lorenz system [30]. For the initial three models, we elucidated the mechanism of TL using a perturbative method and an averaging approximation. In contrast, the coupled Lorenz system necessitated a self-consistent approach instead of the averaging method. In this paper, we extend our previous results for the coupled Lorenz system and provide detailed calculations. We present results not only for the coupled Lorenz system [30] which is based on the original Lorenz system [31], but also for coupled systems constructed from the generalized Lorenz system [32, 33] and the hyperchaotic Lorenz system [34]. The generalized Lorenz system, an extension of the original Lorenz system incorporating the influence of a magnetic field [32], has been shown to exhibit intriguing intermittent dynamics [33]. The hyperchaotic Lorenz system [34], also an extension of the original Lorenz system, displays more complex dynamics than the original Lorenz system [31] and is of particular interest for applications in secure communication and cryptography.
@@ II.3 Numerical simulations
where $\mathrm{CV}$ denotes the coefficient of variation. Namely, $\chi$ is the CV of the mean field of $z_{i}(t)$ normalized by the maximum CV among the individual $z_{i}(t)$. With this definition, $0\leq\chi\leq 1$, where $\chi\simeq 0$ indicates an unsynchronized state and $\chi\simeq 1$ signifies a well-synchronized state. The quantity $\zeta_{i}(t)$ represents the ratio between time series; it becomes approximately constant when WP holds. While the choice of $j$ is arbitrary, we select $j=50$, assuming the parameters $\mu_{i}$ satisfy $\mu_{1}<\mu_{2}<\cdots<\mu_{100}$. By arranging them in this manner, we can expect $z_{50}(t)\simeq z_{0}(t)$. Unless otherwise noted, $\mu_{i}$ is selected from a uniform distribution between $-1.5$ and $1.5$.
@@
Next, we examine the dependence of the TL parameters on the coupling strength. According to the analytical derivation, the numerical simulation results are expected to agree well with the theoretical predictions as $D_{x}$ increases. We verify this expectation using the coupled Lorenz system [30]; the results are shown in Fig. 2. Simulations are performed up to $t=3500$, and TL is computed using the time series from $t=3000$ to $t=3500$. The error bars represent the standard deviation of ten calculations with different initial conditions and realizations of $\rho_{i}$. Initial conditions for numerical simulations are randomly chosen from a uniform distribution between 0 and 1. Depending on the coupling strength, oscillation quenching may occur, rendering it impossible to define $\chi$. Thus, we judge that oscillation quenching occurs when $(1/N)\sum_{i=1}^{N}\langle\left(z_{i}-\langle z_{i}\rangle_{t}\right)^{2}\rangle_{t}$ falls below a certain threshold value; such cases are excluded. We can confirm that qualitatively the same results are obtained for several different threshold values. This procedure is applied to all systems described later. Figure 2 illustrates that when the coupling strength $D_{x}$ becomes comparable to $\max\{\mu_{i}\}$, $\chi$ begins to rise, indicating the onset of synchronization. With a further increase in coupling strength, $\chi\simeq 1$ and $R^{2}\simeq 1$, and the theoretical predictions and numerical simulation results for both $\log\alpha_{\rm t,s}$ and $\beta_{\rm t,s}$ show excellent agreement.
Figure 2: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled Lorenz system. $D_{y}=D_{z}=0,N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
@@ III.1 Model
where $x(t)$, $y(t)$, and $z(t)$ are the original variables of the Lorenz system, and $\sigma$, $\rho$, and $b$ are its original parameters. The variable $w(t)$ and parameters $a$, $c$, $k$, and $l$ are newly introduced to extend the original Lorenz system. By appropriately choosing the parameters $a$, $c$, $k$, and $l$, we can recover the original/generalized/hyperchaotic Lorenz system; when $a=c=k=l=0$, the system reduces to the original Lorenz system (2); when $a=-\omega_{0}$, $c=\omega_{0}$, $k=0$, and $l=\sigma_{\rm m}$, it corresponds to the generalized Lorenz system (24); and when $a=1$, $c=0$, $k=1$, and $l=-\gamma$, it corresponds to the hyperchaotic Lorenz system (25). We refer to the system described by Eq. (26) as the extended Lorenz system. Following the formulation of the coupled Lorenz system used in Ref. [30], we construct the following coupled system for the extended Lorenz system:
@@ III.2 Analytical derivation
When $|\sigma(y_{0}-x_{0})|,|aw_{0}|\gg|\mu_{i}p|,|{\varepsilon}_{i}\sigma q|,|\mu_{i}\tau\sigma(y_{0}-x_{0})|,|\mu_{i}\tau aw_{0}|$, the following equations can be derived by applying the same procedure as before.
@@ III.3 Numerical simulations
Figure 3: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled generalized Lorenz system. $a=-1,c=1,k=0,l=20$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Figure 4: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
@@
Next, we investigate the effects of the other coupling coefficients, $D_{y},D_{z},$ and $D_{w}$, on the validity of the theory. Based on the analytical derivation, we expect that the theoretical framework to become invalid when $D_{y}$ or $D_{z}$ is non-negligible. In contrast, our theory predicts that increasing $D_{w}$ does not affect the results. To test these expectations, we perform numerical simulations using the coupled hyperchaotic Lorenz system (Figs. 6, 7, and 8).
@@
Figure 8: Dependence of TL parameters on $D_{w}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{x}=1000,D_{y}=D_{z}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponhents of spatial TL. (f) Intercepts of spatial TL.
@@ IV Discussion
In this study, building upon our previous work [15], we investigated the emergence of WP and the associated TL with an exponent 2 in coupled Lorenz systems. In particular, we extended the results obtained in Ref. [15] by considering a coupled system constructed from the extended Lorenz system (26), which is newly introduced in this study. We verified our analytical derivation through numerical simulations using coupled systems constructed from the original, generalized, and hyperchaotic Lorenz systems. In the case of the four-variable model, we numerically observe that $|aw_{0}|$ is not always large and our assumption of large $|aw_{0}|$ in our perturbative approach might be violated. Nevertheless, the numerical simulation results are still in good agreement with the theoretical predictions. This observation suggests that the treatment of the $aw_{0}$ term in our perturbative approach requires further consideration.